9. Price-discriminating monopolist Sam owns a plot of land in the desert that isnt worth much. One day, a giant meteorite falls on his property, making a large crater. The event attracts scientists and tourists, and Sam decides to sell nontransferable admission tickets to the meteor crater to both types of visitors: scientists (Market A) and tourists (Market B). The following graphs show daily demand (D) curves and marginal revenue (MR) curves for the two markets. Sams marginal cost of providing admission tickets is zero Market A Market B 10 10 MR MR D 0 2 46 8 10 12 14 16 18 20 QUANTITY (Admission tickets) 0 2 4 68 10 12 14 16 18 20 QUANTITY (Admission tickets)
Answer #1

At price $4, total demand is 16 tickets
(At price $4, in market A demand is 12 tickets and in market B demand is 4 tickets so total market demand is 16 tickets)

$5 in market A and $3 in market B. He will sell a total quantity of 16 tickets
(Monopolist maximises revenue by setting MR = MC and it is given that MC = 0. So, MR = MC = 0. This determines price $5 and quantity 10 tickets in market A. In market B, MC = MR = 0 rule determines price = $3 and quantity = 6 tickets. So, total quantity sold = 10 tickets + 6 tickets = 16 tickets)

Total revenue
Nondiscriminatory = price*total market demand = 4*16 = 64
Discriminatory = (5*10) + (3*6) = 50 + 18 = 68
(As in market A, total revenue = price charged in market A*quantity sold in market A = 5*10 = 50 and in market B, total revenue =  price charged in market B*quantity sold in market B = 3*6 = 18. So, Total revenue = 50 + 18 = 68)

Sam charges a higher price in the market with a relatively low price elasticity of demand.